OpenStudy (anonymous):

A number cube with faces numbered 1 through 6 is rolled and a coin is tossed. What is the probability that the number cube will show a number greater than 4 and the coin will land with tails up?

OpenStudy (tbnrfrags):

2/6 1/2

OpenStudy (anonymous):

The probability of a number greater than 4 (i.e. 5 or 6) is \[\frac{2}{6}\] and the probability of a tails is \[\frac{1}{2}\] so the probability of greater than 4 and tails is: \[\frac{2}{6}*\frac{1}{2}=\frac{2}{12}=\frac{1}{6}\] or about 16.66%

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You can check this by listing every possibility and then counting the ones that meet the conditions: 1H 2H 3H 4H 5H 6H 1T 2T 3T 4T 5T 6T That's 12 possibilities. The ones that meet our conditions are: 5T 6T or 2 of them so: \[\frac{2}{12}=\frac{1}{6}\] Same as before.

OpenStudy (anonymous):

ok thanks. i think i have one more question

OpenStudy (anonymous):

Ok